A company produces 300 units of a product a day, which costs $12 per unit. It employs 3 workers who work 8 hours a day at an hourly rate of $12 dollars per hour. The cost for materials is $1,200 and overhead is $260. The company plans to outsource some of its work to the outside suppliers. The administration believes it will allow lay off one of the workers (use only two workers), pay the other two workers an hourly rate of $15 per hour and produce 350 units per day. But due to the changes the cost of materials will become $1,400 and overhead will become $310.

a. Calculate labor productivity in units/per hour, for both systems of production and compare the results.

b. Calculate labor productivity in units/per dollar, for both systems of production and compare the results.

c. Calculate multifactor productivity in dollar/per dollar for each system of operations (labor + materials + overhead) and compare the results.

d. Calculate the percentage of growth/decline in the multifactor productivity with the new system. Would you recommend the company to outsource this part of the operation to the supplier? Why?
Answer #1

Daily expenses in case I -

Labour cost = 3x8x12 = 288

Material cost = 1200

Overhead = 260

Total =1748

Daily expenses in case II

Labour cost = 2x8x15 =240

Material Cost =1400

Labour cost =310

Total =1950

(a) Productivity in case I = 300 / 3x8 = 12.5

Productivity in case II = 350 /2x8 = 21.87

(b) Productivity in case I = 300 /3x8x12 =1.04

Productivity in case II = 350 /2x8x15 = 1.45

(c) Multifactor Productivity in case I = 300 units /1748 dollars = 0.171 unit /dollar

Multifactor Productivity in case II = 350 units / 1950 dollars = 0.1794 units /dollar

MFP in dollars / dollar = dollar earned / dollars spent in material + labour + overhead

Case I = 300x12 /1748 =2.052

Case II = 350x12 /1950 = 2.152

(d) Percentage growth = MFP in case II / MFP n case I

= 2.152 /2.052 = 1.0487

The growth is 4.87%.

The outsourcing decision is not advisable as the business is saving on its operations.